40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 20 21:29:41 2018, Jon, Omnistructure, Upgrade, Vacuum Controls Switchover Completed 40m_vacuum_acromag_channels.pdf
    Reply  Fri Dec 21 11:13:13 2018, gautam, Omnistructure, VAC, N2 line valved off VacGauges.pngScreenshot_from_2018-12-21_13-02-06.png
       Reply  Fri Dec 21 12:57:10 2018, Koji, Omnistructure, VAC, N2 line valved off 
          Reply  Thu Jan 3 15:08:37 2019, gautam, Omnistructure, VAC, Vac status unknown Screenshot_from_2019-01-03_15-19-51.pngScreenshot_from_2019-01-03_15-14-14.png997B13A9-CAAF-409C-A6C2-00414D30A141.jpeg
             Reply  Fri Jan 4 17:43:24 2019, gautam, Update, CDS, Timing issues 
                Reply  Wed Jan 9 11:33:35 2019, gautam, Update, CDS, Timing issues still persist gpstimeSync.pngScreenshot_from_2019-01-09_17-56-58.png
          Reply  Fri Jan 4 10:25:19 2019, Jon, Omnistructure, VAC, N2 line valved off 
    Reply  Thu Mar 21 18:36:59 2019, Jon, Omnistructure, Upgrade, Vacuum Controls Switchover Completed 40m_Vacuum_Acromag_Channels_20190321.pdf
Message ID: 14375     Entry time: Thu Dec 20 21:29:41 2018     Reply to this: 14377   14493
Author: Jon 
Type: Omnistructure 
Category: Upgrade 
Subject: Vacuum Controls Switchover Completed 

[Jon, Chub, Koji, Gautam]

Summary

Today we carried out the first pumpdown with the new vacuum controls system in place. It performed well. The only problem encountered was with software interlocks spuriously closing valves as the Pirani gauges crossed 1E-4 torr. At that point their readback changes from a number to "L OE-04, " which the system interpreted as a gauge failure instead of "<1E-4." This posed no danger and was fixed on the spot. The main volume was pumped to ~10 torr using roughing pumps 1 and 3. We were limited only by time, as we didn't get started pumping the main volume until after 1pm. The three turbo pumps were also run and tested in parallel, but were isolated to the pumpspool volume. At the end of the day, we closed every pneumatic valve and shut down all five pumps. The main volume is sealed off at ~10 torr, and the pumpspool volume is at ~1e-6 torr. We are leaving the system parked in this state for the holidays. 

Main Volume Pumpdown Procedure

In pumping down the main volume, we carried out the following procedure.

  1. Initially: All valves closed (including manual valves RV1 and VV1); all pumps OFF.
  2. Manually connected roughing pump line to pumpspool via KF joint.
  3. Turned ON RP1 and RP2.
  4. Waited until roughing pump line pressure (PRP) < 0.5 torr.
  5. Opened V3.
  6. Waited until roughing pump line pressure (PRP) < 0.5 torr.
  7. Manually opened RV1 throttling valve to main volume until pumpdown rate reached ~3 torr/min (~3 hours on roughing pumps).
  8. Waited until main volume pressure (P1a/P1b) < 0.5 torr.

We didn't quite reach the end of step 8 by the time we had to stop. The next step would be to valve out the roughing pumps and to valve in the turbo pumps.

Hardware & Channel Assignments

All of the new hardware is now permanently installed in the vacuum rack. This includes the SuperMicro rack server (c1vac), the IOLAN serial device server, a vacuum subnet switch, and the Acromag chassis. Every valve/pump signal cable that formerly connected to the VME bus through terminal blocks has been refitted with a D-sub connector and screwed directly onto feedthroughs on the Acromag chassis.

The attached pdf contains the master list of assigned Acromag channels and their wiring.

Attachment 1: 40m_vacuum_acromag_channels.pdf  72 kB  Uploaded Thu Dec 20 22:36:00 2018  | Hide | Hide all
40m_vacuum_acromag_channels.pdf 40m_vacuum_acromag_channels.pdf 40m_vacuum_acromag_channels.pdf
ELOG V3.1.3-